whitemonsterbunny17 (whitemonsterbunny17):

¢αη ѕσмєσηє ρℓєαѕє ¢нє¢к му αηѕωєяѕ? тнαηкѕ! (ノ^ヮ^)ノ*:・゚✧*:・゚✧

whitemonsterbunny17 (whitemonsterbunny17):

8.3 x 10^6 kg / 2.5 x 10^9 cm^3 3.32 x 10^15 ?

whitemonsterbunny17 (whitemonsterbunny17):

Is that right?

OpenStudy (anonymous):

woah

OpenStudy (anonymous):

10^15?

OpenStudy (anonymous):

You can't convert cm^3 to kg...

whitemonsterbunny17 (whitemonsterbunny17):

It's not 10^15?

whitemonsterbunny17 (whitemonsterbunny17):

What do I do with the exponents then?

OpenStudy (anonymous):

Ok lets take this one step at a time :P

whitemonsterbunny17 (whitemonsterbunny17):

Okay. ^.^

OpenStudy (anonymous):

The units stay

whitemonsterbunny17 (whitemonsterbunny17):

Okay.

OpenStudy (anonymous):

So what's 8.3/ 2.5?

whitemonsterbunny17 (whitemonsterbunny17):

I got 3.32

OpenStudy (anonymous):

That's right now, look at the exponents we have \[\frac{ 10^6 }{ 10^9 }\] what do we do when we divide exponents?

whitemonsterbunny17 (whitemonsterbunny17):

Subtract?

OpenStudy (anonymous):

Yeup ^.^ \[\frac{ x^m }{ x^n } = x^{m-n}\]

whitemonsterbunny17 (whitemonsterbunny17):

So -3?

OpenStudy (anonymous):

Too good

whitemonsterbunny17 (whitemonsterbunny17):

Oh... so it's 3.32 x 10^(-3) ?

OpenStudy (anonymous):

yeahhhhhh!

whitemonsterbunny17 (whitemonsterbunny17):

YAYYYY!!!! :D

whitemonsterbunny17 (whitemonsterbunny17):

\(\Large\bf\rlap{\color{red}{\bigstar~T}\color{gold}{H}\color{greenyellow}{A}\color{lime}{N}\color{cyan}{K~}\color{blue}{Y}\hspace{0 pt}\color{purple}{O}\color{magenta}{U}\color{orchid}{!~\bigstar}}{\color{black}{\; \bigstar~THANK~YOU!~\bigstar}}\)